# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 4.2.2. Descent results
One of the key technical components of showing Theorem 1.13 (2) is the following descent result.
@@
Thanks to Daniel’s work [19] on introducing the category of variations of loop Hodge structures [19], we are able to study harmonic bundles using the classical tools of Hodge theory, notably the existence of a period map, which record the information of the structure of the induced Higgs sheaf as in the following statement.
@@
Let $(E,D,h)$ be a harmonic bundle over a simply connected complex manifold $X$ and $f:M\rightarrow\mathcal{D}$ be the induced period map introduced in [19, Section 3.2]. Write $D=\partial+\bar{\partial}+\theta+\theta^{*}$ by the decomposition (4.4) and set $\mathcal{E}_{X}:=(E,\bar{\partial})$, then there exists a locally free coherent sheaf $\mathcal{E}_{D}$ on $\mathcal{D}$ such that $\mathcal{E}_{X}:=f^{*}\mathcal{E}_{D}$.
To see this, it’s necessary to outline Daniel’s construction of the period map. The reader is encouraged to have a look on [19] for a concise definition when reading the following construction because involving concepts not present in the classic case.
@@
Let $x_{0}\in x$ be a base point and $\mathcal{H}=L^{2}(S^{1},\operatorname{End}(E))$, then the fiber of the canonical outgoing Krein bundle $(\mathcal{H},\mathcal{B},\mathcal{T})$ induced by $(E,h)$ over $x_{0}$ can be identified with the canonical Krein space $(H:=L^{2}(S^{1},\mathbb{C}^{r}),B,T)$ by the identification of $(E,h)_{x_{0}}$ with the standard Hermitian inner product on $\mathbb{C}^{r}$ (see [19, Section 2.1.3]), where $r=\operatorname{rank}(E)$ and $S^{1}=\{z\in\mathbb{C}:|z|=1\}$.
When there is no risk of confusion, the trivial bundle $Y\times H$ over a complex manifold $Y$ will be simply denoted by $H_{Y}$. Set
@@
for any $\lambda\in S^{1}$. Then $\{D_{\lambda}\}$ can be viewed a flat connection $\widetilde{D}$ on $\mathcal{H}$; moreover, $\mathcal{B}$ and $\mathcal{T}$ are $\widetilde{D}-$flat (see [19, Appendix A]). In particular, the canonical Krein space with flat connection $\widetilde{D}$ has the trivialization
$(\mathcal{H},\mathcal{B},\mathcal{T},\widetilde{D})\cong(H_{X},B_{X},T_{X},d_{X}),$
where $B_{X}$ (resp. $T_{X}$) is the trivial Krein metric (resp. the trivial outgoing operator) by identifying with the base point and $d_{X}$ is the tribial connection.
Using Fourier series decomposition, any element $f\in\mathcal{H}$ can be written as $s(\lambda)=\sum\limits_{n\in\mathbb{Z}}s_{n}\lambda^{n}$ such that $(s_{n}(x))_{n\in\mathbb{Z}}\in l^{2}(C^{r})$ for any $x\in X$. The canonical outgoing subbundle is given by $\mathcal{W}=L^{2}_{+}(S^{1},\operatorname{End}(E))$ (cf. [19, Section 2.1.3]), the set of all Fourier series with non-negative degrees. In particular, for any $x$, $\mathcal{W}_{x}$ can be identified with an outgoing subspace $W$ of $H$, which varies when $x$ varies over $X$. In summary, we have the following isomorphism of variations of loop Hodge structures
(4.11) $\gamma:(\mathcal{H},\mathcal{B},\mathcal{T},\mathcal{W},\widetilde{D})\cong(H_{X},B_{X},T_{X},W_{X},d_{X}),$
where $(H_{X},B_{X},T_{X})$ is trivial with respect to $d_{X}$ while $W$ is not.
Denote $\mathrm{Gr}(H)$ be the outgoing Grassmannian of $H$, that is the set of all outgoing subspaces of $H$. By [19, Proposition 3.7 and Remark 3.8], we can identify $\mathrm{Gr}(H)$ with $\Lambda_{\sigma}^{\infty}G/K$ (cf. [19, p. 619] for its definition), where $G=\mathrm{GL}(r,\mathbb{C})$ and $K=\mathrm{U}(r,C)$. Therefore, $f:x\mapsto W_{x}$ gives a map from $X$ to $\Lambda_{\sigma}^{\infty}G/K$. Denote $\mathcal{D}=\Lambda_{\sigma}G/K$ be the period domain defined in [19, Section 3.1.4]. $\mathcal{D}$ is an infinite-dimensional complex Hilbert manifold and $f(x)$ indeed lies in $\mathcal{D}$ (cf. [19, Lemma 3.19]). Then $f:X\rightarrow\mathcal{D}$ is defined as the period map of $(E,D,h)$.
@@
(4.12) $(E,\bar{\partial})\cong(\mathcal{W}/\mathcal{T}\mathcal{W},\widetilde{D}^{0,1})\cong(W_{X}/T_{X}W_{X},d^{0,1})$
@@
$\mathrm{Gr}(H)=\Lambda_{\sigma}^{\infty}G/K$ implies that there exists a tautological bundle over $\mathcal{D}$ defined by $(H^{0,\mathcal{D}})_{W\in\mathcal{D}}$, where $H^{0,\mathcal{D}}$ is the orthogonal complement of $TW$ in $W$ (cf. [19, Definition 3.12]). Let $d_{\mathcal{D}}$ be the trivial connection of the trivial bundle $H_{\mathcal{D}}$, then it induces a holomorphic structure $\bar{\partial}_{\mathcal{D}}$ on $H^{0,\mathcal{D}}$. According to the construction of the period map $f$ explained as above,
$(W_{X}/T_{X}W_{X},d_{X}^{0,1})=f^{*}(H^{0,D},d_{\mathcal{D}}^{0,1})$
since $f$ is holomorphic by [19, Theorem 3.18]. Via the identification (4.12), we conclude that $\mathcal{E}_{X}=(E,\bar{\partial})=f^{*}(H^{0,D},d_{\mathcal{D}}^{0,1})$. Taking $\mathcal{E}_{\mathcal{D}}=(H^{0,D},d_{\mathcal{D}}^{0,1})$ and the proof is complete. ∎
Building on Lemma 4.12 and [19, Theorem 1.2], we can adapt the arguments of [37, Corollary 5.8] to prove Proposition 4.11, where treats Higgs bundles from variations of Hodge structures.
@@
Recall Takayama’s result [101] that for any $x\in X$, there exists an open neighborhood $U$ of $x$ such that $M:=\pi^{-1}(U)$ is simply connected. We shall show that there exists a locally free sheaf $\mathcal{E}_{U}$ on $U$ such that $\mathcal{E}_{\widehat{X}}|_{\pi^{-1}(U)}=\pi^{*}\mathcal{E}_{U}$, which suffices to conclude the desired statement by Lemma 2.4.
Let $f:\pi^{-1}(U)\rightarrow\mathcal{D}$ be the induced period map of the harmonic bundle $(E,D,h)|_{\pi^{-1}(U)}$. Building on Lemma 4.12, there exists a locally free sheaf $\mathcal{E}_{\mathcal{D}}$ such that $\mathcal{E}_{X}|_{\pi^{-1}(U)}=f^{*}\mathcal{E}_{\mathcal{D}}$. If $f$ factories through $\pi$, i.e., $f=g\circ\pi$ for some holomorphic map $g:U\rightarrow\mathcal{D}$, then $\mathcal{E}_{X}=\pi^{*}(g^{*}\mathcal{E}_{\mathcal{D}})$. It suffices to show that $f$ is constant on $\pi^{-1}(x)$ for any $x\in U$. Fix a point $x\in U$. For any rational curve $\gamma:\mathbb{P}^{1}\rightarrow\pi^{-1}(x)$, $f\circ\gamma$ is again holomorphic and horizontal (cf the definition in [19, 3.1.6]) and so is a period map induced by a harmonic bundle structure over $\mathbb{P}^{1}$ (cf. [19, Theorem 1.2]). However, any harmonic bundle structure over $\mathbb{P}^{1}$ is a trivial bundle and so $x\mapsto W_{x}$ in (4.11) is constant, which implies that $f\circ\gamma$ is a constant map. Recall that $\pi^{-1}(x)$ is rationally chain connected (cf. [30]), i.e., any two points of $\pi^{-1}(x)$ can be connected by a chain of rational curves. We conclude that $f$ is constant on $\pi^{-1}(x)$ for any $x\in U$ and the proof is complete. ∎
@@
Let us complete the proof of Theorem 1.13. Let $X$ be a compact Kähler klt space and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{pHiggs}_{X_{\mathrm{reg}}}$. Building on Theorem 3.2, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is locally free and admits a pluri-harmonic metric $H_{X\setminus Z}$ outside an analytic subspace $Z$ such that $\mathrm{codim}_{X}Z\geq 2$ and $X\setminus Z$ is smooth. It follows from of Proposition 4.8 that the associated harmonic bundle structure of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus Z}$ extends to a harmonic bundle on $X_{\mathrm{reg}}$ and so $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus Z}$ extends to a locally free Higgs sheaf, which coincides with $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ by the reflexivity.
Thus we have verified the “only if part” of Theorem 1.13 (2); moreover, Theorem 1.13 (1) directly follows from Proposition 4.8 by substituting $X\setminus Z=X_{\mathrm{reg}}$. The remaining statements of Theorem 1.13 are included by the following statement.
@@
Let $X$ be a compact Kähler klt space space. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a locally free Higgs sheaf equipped with a pluri-harmonic metric $H_{\mathrm{reg}}$ on $X_{\mathrm{reg}}$. Then for any maximally quasi-étale cover $\gamma:Y\rightarrow X$ induced by some $G\subset\mathrm{Aut}(Y)$, the following hold:
  1. ($4.1$)
    $\gamma^{[*]}\mathcal{E}_{X}$ is locally free and in particular $\gamma^{[*]}\theta_{{X_{\mathrm{reg}}}}$ can be extended to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}=\gamma^{[*]}\mathcal{E}_{X}$.
  2. (4.14.2)
@@
(4.13) $\begin{split}&\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{F}_{Y})=\deg_{(\pi^{*}\omega_{0},\cdots,\pi^{*}\omega_{n-2})}(\mathcal{F}_{\widehat{Y}})\\ &=-\int_{\widehat{Y}\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}}\sqrt{-1}\operatorname{tr}(\partial\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}}\wedge\bar{\partial}\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}})\wedge\pi^{*}\omega_{0}\wedge\cdots\pi^{*}\omega_{n-2}\\ &\leq 0=\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{Y})\end{split}$
@@
since $F_{H_{\widehat{Y}},\theta_{\widehat{Y}}}=0$ and $c_{1}(\mathcal{E}_{Y})=0$ by (4.1.1), where $\Sigma_{\mathcal{F}_{\widehat{Y}}}$ is the singular set of $\mathcal{F}_{\widehat{Y}}$. Suppose that $(\mathcal{E}_{Y},\theta_{Y})$ is not $(\omega_{0},\cdots,\omega_{n-2})$-stable, then there exists some $\mathcal{F}_{Y}$ such that “=” in the last line of (4.13) holds, then $\partial\pi_{\widehat{Y}}^{H_{\widehat{Y}}}=0$ on $\pi^{-1}(Y_{\mathrm{reg}})\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}$ since $\pi^{*}\omega_{0},\cdots,\pi^{*}\omega_{n-2}$ are strictly positive on $\pi^{-1}(Y_{\mathrm{reg}})$. Thus
$(\mathcal{E}_{\widehat{Y}},\theta_{\widehat{Y}})=(\mathcal{F}_{\widehat{Y}},\theta_{\widehat{\mathcal{F}}})\oplus(\mathcal{F}_{\widehat{Y}}^{\perp},\theta_{\widehat{\mathcal{F}}^{\perp}})$
@@
splits as locally free Higgs sheaves on $\pi^{-1}(Y_{\mathrm{reg}})\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}$, where $\theta_{\widehat{\mathcal{F}}}$ (resp. $\theta_{\widehat{\mathcal{F}}^{\perp}}$) is the induced Higgs field on $\mathcal{F}_{\widehat{Y}}$ (resp. $\mathcal{F}_{\widehat{Y}}^{\perp}$). Consequently,
$(\mathcal{E}_{{Y}},\theta_{{Y}})=(\mathcal{F}_{{Y}},\theta_{{\mathcal{F}}})\oplus(\mathcal{F}_{{Y}}^{\perp},\theta_{{\mathcal{F}}^{\perp}})$
@@
splits as locally free Higgs sheaves on $Y_{\mathrm{reg}}\setminus\Sigma_{\mathcal{F}_{Y}}$, where $\mathcal{F}_{Y}^{\perp}$ is the orthogonal complement with respect to $\pi_{*}H_{\widehat{Y}}$ and $\Sigma_{\mathcal{F}_{Y}}$ is the singular set of $\mathcal{F}_{Y}$. Then the induced Hermitian metric $H_{\mathcal{F}_{Y}}$ (resp. $H_{\mathcal{F}_{Y}^{\perp}}$) on $\mathcal{F}_{Y}$ (resp. $\mathcal{F}_{Y}^{\perp}$) is pluri-harmonic. Since $\Sigma_{\mathcal{F}_{Y}}$ has codimension at least $2$, then $(\mathcal{F}_{Y},\theta_{\mathcal{F}_{Y}})$ and $(\mathcal{F}_{Y}^{\perp},\theta_{\mathcal{F}_{Y}^{\perp}})$ extend to a locally free Higgs sheaf that arises from a harmonic bundle structure by the same argument as above. Applying an induction argument, we verify the $(\omega_{0},\cdots,\omega_{n-2})$-polystability of $(\mathcal{E}_{Y},\theta_{Y})$.
@@
Let us complete the proof. The item (4.14.2), the item (4.1.1), Remark 2.22 and Lemma 2.21 together implies that the orbifold Chern classes of $\mathcal{E}_{X}$ of $\mathcal{E}_{X_{\mathrm{reg}}}$ vanishes. On the other hand, the item ((4.14.3)) implies $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is stable with respect to $\gamma^{*}\omega_{0},\cdots,\gamma^{*}\omega_{n-2}$ for any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$ and in particular, is $G$-stable, which implies the $(\omega_{0},\cdots,\omega_{n-2})$-polystability of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ by Lemma 2.24. ∎
5. Non-abelian Hodge correspondence over Kähler klt spaces

### v2

@@ 4.2.2. Descent results
A key technical step in proving Theorem 1.13 (2) is the following descent result.
@@
Thanks to Daniel’s work on variations of loop Hodge structures [19], we are able to study harmonic bundles using the classical tools of Hodge theory, in particular the existence of a period map, which encodes the structure of the induced Higgs sheaf as follows.
@@
Let $(E,D,H)$ be a harmonic bundle over a simply connected complex manifold $X$ and $f:X\rightarrow\mathcal{D}$ be the induced period map introduced in [19, Section 3.2]. Write $D=\partial+\bar{\partial}+\theta+\theta^{*}$ according to the decomposition (4.4) and set $\mathcal{E}_{X}:=(E,\bar{\partial})$. Then there exists a locally free coherent sheaf $\mathcal{E}_{D}$ on $\mathcal{D}$ such that $\mathcal{E}_{X}:=f^{*}\mathcal{E}_{D}$.
To see this, we briefly outline Daniel’s construction of the period map. For a concise formulation, we refer to [19]; below we summarize the essential ideas.
@@
Fix a base point $x_{0}\in X$ and let $\mathcal{H}=L^{2}(S^{1},\operatorname{End}(E))$. Then the fiber at $x_{0}$ of the canonical outgoing Krein bundle $(\mathcal{H},\mathcal{B},\mathcal{T})$ induced by $(E,H)$ can be identified with the canonical Krein space $\widehat{H}:=(L^{2}(S^{1},\mathbb{C}^{r}),B,T)$ via the identification of $(E,H)_{x_{0}}$ with the standard Hermitian inner product on $\mathbb{C}^{r}$ (see [19, Section 2.1.3]), where $r=\operatorname{rank}(E)$ and $S^{1}=\{z\in\mathbb{C}:|z|=1\}$.
When there is no risk of confusion, the trivial bundle $Y\times\widehat{H}$ over a complex manifold $Y$ will be denoted simply by $\widehat{H}_{Y}$. Set
@@
for any $\lambda\in S^{1}$. Then $\{D_{\lambda}\}$ can be regarded as a flat connection $\widetilde{D}$ on $\mathcal{H}$; moreover, $\mathcal{B}$ and $\mathcal{T}$ are $\widetilde{D}-$flat (see [19, Appendix A]). In particular, the canonical Krein space with the flat connection $\widetilde{D}$ admits a trivialization
$(\mathcal{H},\mathcal{B},\mathcal{T},\widetilde{D})\cong(\widehat{H}_{X},B_{X},T_{X},d_{X}),$
where $B_{X}$ (resp. $T_{X}$) is the trivial Krein metric (resp. trivial outgoing operator) by identification with the base point, and $d_{X}$ is the trivial connection.
Via Fourier series expansion, any element $s\in\mathcal{H}$ can be written as $s(\lambda)=\sum\limits_{n\in\mathbb{Z}}s_{n}\lambda^{n}$ with $(s_{n}(x))_{n\in\mathbb{Z}}\in l^{2}({\mathbb{C}}^{r})$ for each $x\in X$. The canonical outgoing subbundle is given by $\mathcal{W}=L^{2}_{+}(S^{1},\operatorname{End}(E))$ (cf. [19, Section 2.1.3]), consisting of all Fourier series with nonnegative degrees. In particular, for any $x$, $\mathcal{W}_{x}$ can be identified with an outgoing subspace $W$ of $\widehat{H}$, and this subspace varies with $x\in X$. In summary, we obtain an isomorphism of variations of loop Hodge structures
(4.11) $\gamma:\left((\mathcal{H},\mathcal{B},\mathcal{T}),\mathcal{W},\widetilde{D}\right)\cong(\widehat{H}_{X},W_{X},d_{X}),$
where $\widehat{H}_{X}=\widehat{H}\times X$ is trivial with respect to $d_{X}$ while $W_{X}$ is not.
[…]
@@
(4.12) $(E,\bar{\partial})\cong(\mathcal{W}/\mathcal{T}\mathcal{W},\widetilde{D}^{0,1})\cong(W_{X}/T_{X}W_{X},d_{X}^{0,1}),$
@@
The identification $\mathrm{Gr}(\widehat{H})=\Lambda_{\sigma}^{\infty}G/K$ implies that there exists a tautological bundle over $\mathcal{D}$, defined by $(\widehat{H}^{0,\mathcal{D}})_{W\in\mathcal{D}}$, where $\widehat{H}^{0,\mathcal{D}}$ is the orthogonal complement of $TW$ in $W$ (cf. [19, Definition 3.12]). Let $d_{\mathcal{D}}$ be the trivial connection on the trivial bundle $\widehat{H}_{\mathcal{D}}$, then it induces a holomorphic structure $\bar{\partial}_{\mathcal{D}}$ on $\widehat{H}^{0,\mathcal{D}}$. According to the construction of the period map $f$ explained as above, we have
$(W_{X}/T_{X}W_{X},d_{X}^{0,1})=f^{*}(\widehat{H}^{0,D},d_{\mathcal{D}}^{0,1})$
since $f$ is holomorphic by [19, Theorem 3.18]. Via the identification (4.12), we conclude that $\mathcal{E}_{X}=(E,\bar{\partial})=f^{*}(\widehat{H}^{0,D},d_{\mathcal{D}}^{0,1})$. Taking $\mathcal{E}_{\mathcal{D}}=(\widehat{H}^{0,D},d_{\mathcal{D}}^{0,1})$, the proof is complete. ∎
In view of Lemma 4.12 and [19, Theorem 1.2], we can adapt the arguments in [37, Corollary 5.8]—which treats Higgs bundles from variations of Hodge structures—to prove Proposition 4.11.
@@
Recall Takayama’s result [104] that for any $x\in X$, there exists an open neighborhood $U$ of $x$ such that $\pi^{-1}(U)$ is simply connected. We shall show that there exists a locally free sheaf $\mathcal{E}_{U}$ on $U$ such that $\mathcal{E}_{\widehat{X}}|_{\pi^{-1}(U)}=\pi^{*}\mathcal{E}_{U}$, which suffices to conclude the desired statement by Lemma 2.4.
Let $f:\pi^{-1}(U)\rightarrow\mathcal{D}$ be the induced period map of the harmonic bundle $(E,D,h)|_{\pi^{-1}(U)}$. By Lemma 4.12, there exists a locally free sheaf $\mathcal{E}_{\mathcal{D}}$ such that $\mathcal{E}_{X}|_{\pi^{-1}(U)}=f^{*}\mathcal{E}_{\mathcal{D}}$. If $f$ factors through $\pi$, i.e., $f=g\circ\pi$ for some holomorphic map $g:U\rightarrow\mathcal{D}$, then $\mathcal{E}_{X}=\pi^{*}(g^{*}\mathcal{E}_{\mathcal{D}})$. It suffices to show that $f$ is constant on $\pi^{-1}(x)$ for $x\in U$. Fix a point $x\in U$. For any rational curve $\gamma:\mathbb{P}^{1}\rightarrow\pi^{-1}(x)$, $f\circ\gamma$ is again holomorphic and horizontal (cf. [19, Section 3.1.6]) and so is a period map induced by a harmonic bundle structure over $\mathbb{P}^{1}$ (cf. [19, Theorem 1.2]). However, any harmonic bundle over $\mathbb{P}^{1}$ is trivial, and hence its period map $x\mapsto W_{x}$ is constant, which implies that $f\circ\gamma$ is a constant map. Recall that $\pi^{-1}(x)$ is rationally chain connected (cf. [30]), i.e., any two points of $\pi^{-1}(x)$ can be connected by a chain of rational curves. We conclude that $f$ is constant on $\pi^{-1}(x)$ for any $x\in U$ and the proof is complete. ∎
@@
We now complete the proof of Theorem 1.13. Let $X$ be a compact Kähler klt space and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{pHiggs}_{X_{\mathrm{reg}}}$. Building on Theorem 3.2, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is locally free and admits a pluri-harmonic metric $H_{X\setminus Z}$ outside an analytic subspace $Z$ with $\mathrm{codim}_{X}Z\geq 2$ and $X\setminus Z$ smooth. It follows from Proposition 4.8 that the harmonic bundle associated to $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus Z}$ extends to a harmonic bundle on $X_{\mathrm{reg}}$. Consequently, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus Z}$ extends to a locally free Higgs sheaf on $X_{\mathrm{reg}}$, which coincides with $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ by reflexivity.
Thus we have verified the “only if ” part of Theorem 1.13 (2). Moreover, Theorem 1.13 (1) follows directly from Proposition 4.8 by taking $X\setminus Z=X_{\mathrm{reg}}$. The remaining statements of Theorem 1.13 are covered by the following proposition.
@@
Let $X$ be a compact Kähler klt space. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a locally free Higgs sheaf equipped with a pluri-harmonic metric $H_{X_{\mathrm{reg}}}$ on $X_{\mathrm{reg}}$. Then for any maximally quasi-étale cover $\gamma:Y\rightarrow X$ induced by some subgroup $G\subset\mathrm{Aut}(Y)$, the following hold:
  1. (4.14.a)
    $\gamma^{[*]}\mathcal{E}_{X}$ is locally free; in particular, $\gamma^{[*]}\theta_{{X_{\mathrm{reg}}}}$ extends to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}=\gamma^{[*]}\mathcal{E}_{X}$.
  2. (4.14.b)
@@
(4.13) $\begin{split}&\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{F}_{Y})=\deg_{(\pi^{*}\omega_{0},\cdots,\pi^{*}\omega_{n-2})}(\mathcal{F}_{\widehat{Y}})\\ &=-\int_{\widehat{Y}\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}}\sqrt{-1}\operatorname{tr}(\partial\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}}\wedge\bar{\partial}\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}})\wedge\pi^{*}\omega_{0}\wedge\cdots\pi^{*}\omega_{n-2}\\ &\leq 0=\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{Y}),\end{split}$
@@
where we used that $F_{H_{\widehat{Y}},\theta_{\widehat{Y}}}=0$ and $c_{1}(\mathcal{E}_{Y})=0$ by (4.1.a), and $\Sigma_{\mathcal{F}_{\widehat{Y}}}$ denotes the singular set of $\mathcal{F}_{\widehat{Y}}$. Now suppose that $(\mathcal{E}_{Y},\theta_{Y})$ is not $(\omega_{0},\cdots,\omega_{n-2})$-stable. Then there exists some $\mathcal{F}_{Y}$ that achieve equality in the last line of (4.13). Since the forms $\pi^{*}\omega_{0},\cdots,\pi^{*}\omega_{n-2}$ are strictly positive on $\pi^{-1}(Y_{\mathrm{reg}})$, this forces $\partial\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}}=0$ on $\pi^{-1}(Y_{\mathrm{reg}})\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}$. Thus
$(\mathcal{E}_{\widehat{Y}},\theta_{\widehat{Y}})=(\mathcal{F}_{\widehat{Y}},\theta_{\widehat{Y}}|_{\mathcal{F}_{\widehat{Y}}})\oplus(\mathcal{F}_{\widehat{Y}}^{\perp},\theta_{\widehat{Y}}|_{\mathcal{F}_{\widehat{Y}}^{\perp}})$
@@
splits as a direct sum of locally free Higgs sheaves on $\pi^{-1}(Y_{\mathrm{reg}})\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}$, where $\mathcal{F}_{\widehat{Y}}^{\perp}$ is the orthogonal complement with respect to $H_{\widehat{Y}}$. Consequently,
$(\mathcal{E}_{{Y}},\theta_{{Y}})=(\mathcal{F}_{{Y}},\theta_{Y}|_{{\mathcal{F}_{Y}}})\oplus(\mathcal{F}_{{Y}}^{\perp},\theta_{Y}|_{{\mathcal{F}}_{Y}^{\perp}})$
@@
splits as locally free Higgs sheaves on $Y_{\mathrm{reg}}\setminus\Sigma_{\mathcal{F}_{Y}}$, where $\mathcal{F}_{Y}^{\perp}$ is the orthogonal complement with respect to $\pi_{*}H_{\widehat{Y}}$ and $\Sigma_{\mathcal{F}_{Y}}$ is the singular set of $\mathcal{F}_{Y}$. Then the induced Hermitian metric $H_{\mathcal{F}_{Y}}$ (resp. $H_{\mathcal{F}_{Y}^{\perp}}$) on $\mathcal{F}_{Y}$ (resp. $\mathcal{F}_{Y}^{\perp}$) is pluri-harmonic. Since $\Sigma_{\mathcal{F}_{Y}}$ has codimension at least $2$, both $(\mathcal{F}_{Y},\theta_{Y}|_{\mathcal{F}_{Y}})$ and $(\mathcal{F}_{Y}^{\perp},\theta_{Y}|_{\mathcal{F}_{Y}^{\perp}})$ extend to locally free Higgs sheaves arising from a harmonic bundle structure, by the same argument as above. Applying an induction argument, we verify the $(\omega_{0},\cdots,\omega_{n-2})$-polystability of $(\mathcal{E}_{Y},\theta_{Y})$.
@@
Finally, we conclude the proof. Item (4.14.b), together with item (4.1.a), Remark 2.22 and Proposition 2.21 implies that all orbifold Chern classes of $\mathcal{E}_{X}$ vanishes. Moreover, item (4.14.c) implies $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is polystable with respect to $\gamma^{*}\omega_{0},\cdots,\gamma^{*}\omega_{n-2}$ for any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$ and in particular, is $G$-polystable. By Lemma 2.24, this yields the $(\omega_{0},\cdots,\omega_{n-2})$-polystability of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$. ∎
5. Non-abelian Hodge correspondence on Kähler klt spaces